# First Step Towards Matter Power Spectrum Reconstruction with Stage-III Weak Gravitational Lensing Surveys

### v1

@@ 6 Conclusions
In this letter, we have taken a first step towards reconstructing the shape of the matter power spectrum measured with Stage-III WL lensing surveys. Reconstructing the matter power spectrum from different surveys that are sensitive to different redshift ranges and scales can give insight into the physical origins of observed tensions in cosmological parameters. We do not assume a cosmological model for the shape of the matter power spectrum but instead remain agnostic to its shape by assuming a double power law in wavenumber and scale factor (eq. 8). This allows us to directly compare the results of different surveys and quantify the shape of the matter power spectrum preferred by the data. We identify the scales to which the data is most sensitive, and we test different assumptions for the strength of the intrinsic alignment signal. The results highlight the potential of directly constraining the matter power spectrum to explore tensions in current data sets. We find
@@
    Stage-III WL surveys can only poorly constrain the shape and evolution of the matter power spectrum in the low-redshift Universe. Still, they prefer a weaker evolution in the redshift range of 0.5 - 0.1 than predicted by $\Lambda$CDM.
@@
    Including intrinsic alignment of galaxies does not significantly impact the fitting accuracy or best-fit parameters.
More accurate measurement by next-generation WL surveys carried out by Euclid (Euclid Collaboration et al., 2024), Roman (Spergel et al., 2015), and Rubin (LSST Science Collaboration et al., 2009) as well as more accurate modeling will help understand the physical origin of the tensions between different observational cosmological probes.
Appendix A KiDS-1000 COSEBIs predicition
Figure 5: KiDS-1000 COSEBIs measurement (black scatter points) with best-fit $A_{\mathrm{IA}}$ = 1 and 0 double power law (solid blue and curves, respectively) for the first 5 COSEBIs $n$-modes. Each sub-panel corresponds to a combination of two tomographic bins, as indicated in the upper-right corner.
This appendix shows the best-fit prediction for the double power law model to the KiDS-1000 COSEBIs measurements. The scatter points in Fig. 5 are the KiDS-1000 measurements (Asgari et al. 2021), and the red and blue curves are the best-fit double power law fits with $A_{\mathrm{IA}}=0$ and $A_{\mathrm{IA}}=1$, respectively. The subpanels show the measured $E_{n}$ as a function of the COSEBI $n$-modes for the combination of two tomographic bins, as indicated in the upper-right corner of each panel. There is no difference in the fitting accuracy with or without including IA, as seen from the reduced chi-square values in Table 2. The double power law estimate fits the same ranges well as the best-fit $\Lambda$CDM prediction from Asgari et al. (2021). For example, the $\Lambda$CDM and our double power law estimate both provide good fits to most of the data, but they underpredict the signal for the high-$n$ modes in the ($1-5$) tomographic bin combination and across the entire range for the ($2-2$) combination.

### v2

@@ 6 Conclusions
In this paper we have taken a first step toward reconstructing the shape of the matter power spectrum measured with Stage III WL lensing surveys. Reconstructing the matter power spectrum from different surveys that are sensitive to different redshift ranges and scales can give insights into the physical origins of observed tensions in cosmological parameters. We did not assume a cosmological model for the shape of the matter power spectrum, but instead remained agnostic to its shape by assuming a double power law in wavenumber and scale factor (Eq. 8). This allowed us to directly compare the results of different surveys and quantify the shape of the matter power spectrum preferred by the data. We identified the scales to which the data are most sensitive, and we tested different assumptions for the strength of the intrinsic alignment signal. The results highlight the potential of directly constraining the matter power spectrum to explore tensions in current data sets. We find the following:
@@
    Stage III WL surveys can only poorly constrain the shape and evolution of the matter power spectrum in the low-redshift Universe. Even so, they favor a weaker evolution in the redshift range of 0.5 - 0.1 than predicted by $\Lambda$CDM.
@@
    Including an intrinsic alignment of galaxies does not significantly impact the fitting accuracy or the best-fitting parameters.
More accurate measurements by next-generation WL surveys carried out by Euclid (Euclid Collaboration et al., 2024), Roman (Spergel et al., 2015), and Rubin (LSST Science Collaboration et al., 2009) as well as more accurate modeling will help us understand the physical origin of the tensions between different observational cosmological probes.
Appendix A KiDS-1000 COSEBI prediction
This Appendix shows the best-fitting prediction for the double power-law model to the KiDS-1000 COSEBIs measurements. The scatter points in Fig. 5 are the KiDS-1000 measurements (Asgari et al., 2021), and the red and blue curves are the best-fitting double power-law fits with $A_{\mathrm{IA}}=0$ and $A_{\mathrm{IA}}=1$, respectively. The subpanels show the measured $E_{n}$ as a function of the COSEBI $n$-modes for the combination of two tomographic bins, as indicated in the upper-right corner of each panel. There is no difference in the fitting accuracy with or without including IA, as seen from the reduced chi-square values in Table 2. The double power-law estimate fits the same ranges well as the best-fitting $\Lambda$CDM prediction from Asgari et al. (2021). For example, the $\Lambda$CDM and our double power-law estimate both provide good fits to most of the data, but they underpredict the signal for the high-$n$ modes in the ($1-5$) tomographic bin combination and across the entire range for the ($2-2$) combination.
Figure 5: KiDS-1000 COSEBIs measurement (black scatter points) with best-fitting $A_{\mathrm{IA}}$ = 1 and 0 double power law (solid blue and curves, respectively) for the first five COSEBIs $n$-modes. Each subpanel corresponds to a combination of two tomographic bins, as indicated in the upper right corner.